# FloodSQL-Bench: A Retrieval-Augmented Benchmark for Geospatially-Grounded Text-to-SQL

### v1

@@ Abstract
FloodSQL-Bench: A Retrieval-Augmented Benchmark for Geospatially-Grounded Text-to-SQL
Hanzhou Liu,   Kai Yin,   Zhitong Chen,   Chenyue Liu,   Ali Mostafavi Texas A&M University
1 Introduction
Text-to-SQL translates natural language questions into executable SQL queries, enabling intuitive access to relational databases through conversational interfaces Katsogiannis-Meimarakis and Koutrika (2023). General-purpose benchmarks such as Spider Yu et al. (2018) and BIRD Li et al. (2023) have significantly advanced research on this task. However, these benchmarks primarily emphasize single-table queries or simple joins in open-domain settings, and thus fail to capture the challenges of domain-specific, multi-table and geospatial reasoning that are crucial for real-world applications.
Flood management is one such domain, where decision-making depends on integrating different data sources, including geographic features, demographic statistics, and critical infrastructure records Cutter et al. (2012); Wing et al. (2018). Unlike open-domain settings, practitioners must combine information across multiple layers, for example linking census tracts with floodplain boundaries or associating hospitals with historical flood claims. To overcome these challenges, we propose FloodSQL-Bench, a novel benchmark constructed from real-world datasets and designed around progressively complex queries, ranging from single-table lookups to multi-table reasoning involving key-based joins, spatial joins, and their hybrids. This design enables systematic evaluation of Text-to-SQL models on realistic geospatial and socio-economic reasoning tasks essential for flood risk assessment and emergency response.
@@ 2 Related Work
Text-to-SQL research has evolved from single-table parsing to cross-domain and multi-table reasoning. Early datasets such as ATIS and GeoQuery focused on domain-specific grammars, whereas Spider Yu et al. (2018) established a large-scale cross-domain benchmark that has become the de facto standard for evaluating Text-to-SQL systems. Subsequent works, including SParC Yu et al. (2019b) and CoSQL Yu et al. (2019a), extended Spider to multi-turn and conversational settings. More recently, benchmark efforts such as TableBench Wu et al. (2025) target broader table-based reasoning beyond SQL generation. However, all of these benchmarks primarily center on open-domain databases with relatively simple join structures, and rarely address domain-specific, multi-table reasoning that involves heterogeneous relational and spatial data, a key challenge for high-stakes applications such as flood risk management.
@@
Geospatial reasoning introduces unique challenges beyond conventional relational semantics, as queries often involve spatial joins, coordinate transformations, and topology-aware aggregation. Prior work on natural-language interfaces to spatial databases Li et al. (2019); Liu et al. (2025) and spatial question answering Chen et al. (2021) has explored spatial relations such as containment, intersection, and proximity. However, these systems are typically evaluated on synthetic or small-scale datasets and do not address realistic, domain-specific reasoning that integrates heterogeneous spatial and tabular information. In contrast, real-world flood management requires the fusion of diverse geospatial sources—such as physical floodplains, demographic indicators, and critical infrastructure layers, to support actionable decision-making Cutter et al. (2012); Wing et al. (2018). Despite their importance, no existing Text-to-SQL or QA benchmarks systematically incorporate such spatially grounded, multi-table reasoning tasks. FloodSQL-Bench bridges this gap by providing a unified benchmark that combines key-based, spatial, and hybrid joins across real disaster-management datasets, enabling rigorous evaluation of large language models (LLMs) under retrieval-augmented (RAG) settings.
@@
Retrieval-Augmented Generation (RAG) has become a central approach for grounding large language models (LLMs) in traditional NLP tasks like question answering and conversation Lewis et al. (2020); Izacard and Grave (2021); Shuster et al. (2021), with recent work extending retrieval beyond unstructured text to structured sources such as relational databases and tables Ayala and Bechard (2024); Wu et al. (2024). However, existing text-to-SQL benchmarks primarily focus on non-spatial schema and do not evaluate reasoning over geometric data, spatial joins, or multi-layer geospatial infrastructures. Although prior work has explored geospatial question answering using knowledge graphs Li et al. (2025) and GIS tool–driven workflow automation Zhang et al. (2024), neither line of research concentrates on SQL-based spatial joins across multiple tables. Recent advances in table-aware retrieval Zhang et al. (2023); Ziletti and DAmbrosi (2024) further highlight the importance of retrieval granularity for structured reasoning, yet these methods have not been explicitly tested in spatial settings where geometry, topology, and coordinate systems must be considered. To the best of our knowledge, FloodSQL-Bench addresses these gaps by providing the first benchmark designed for multi-layer geospatial SQL reasoning and RAG evaluation, enabling systematic assessment of LLMs across lookup, relational, and spatially grounded analytical queries.
3 Benchmark Construction
@@
To balance realism with tractability, FloodSQL-Bench focuses on three flood-prone states, Texas, Florida, and Louisiana, which together account for a disproportionate share of National Flood Insurance Program (NFIP) claims and Federal Emergency Management Agency (FEMA)–declared disasters. Within this scope, FloodSQL-Bench integrates ten heterogeneous tables across three spatial aspects: (i) non-spatial layers, claims, svi111svi: Social Vulnerability Index., cre222cre: Community Resilience Estimates., nri333nri: National Risk Index.; (ii) polygon layers, floodplain, census_tracts, zcta, county; and (iii) point layers, schools, hospitals. These layers reflect key information needed for flood risk management Cutter et al. (2012); Wing et al. (2018).
@@
3.1 Tabular Data Foundations
@@
FloodSQL-Bench adopts standardized geographic identifiers defined by the U.S. Census Bureau to ensure consistent key-based joins across heterogeneous tables U.S. Census Bureau (2020). At the tract level, an 11-digit GEOID encodes a 2-digit state code, a 3-digit county code, and a 6-digit tract code, serving as the primary join key among the census_tracts, claims, svi, nri, and cre tables. At the county level, the 5-digit prefix of the tract GEOID uniquely identifies each county in the county and hospitals tables. At the ZIP level, the ZIP field serves as the geographic identifier for the schools and hospitals tables. Two spatial layers, zcta and floodplain, lack non-spatial join keys and are instead linked through spatial relationships only.
@@
FloodSQL-Bench standardizes spatial representations across various data sources. Specifically, polygon layers retain geometries to support polygon–polygon joins via spatial SQL functions such as ST_Intersects, ST_Contains, and etc. In contrast, point layers store only explicit latitude and longitude (LAT&LON) fields, with geometries constructed on demand using ST_Point(LON, LAT). To ensure efficiency, all geometries are projected to a common coordinate reference system (CRS) Goodchild (1992); Burrough et al. (2015) and simplified with topology-preserving tolerances Visvalingam and Whyatt (2017); GEOS Development Team .
3.2 Join Semantics and Relations
@@ 3.4 Question and SQL Annotation
We follow the procedure of TableBench Wu et al. (2025) to construct diverse question–SQL pairs. Specifically, we first manually design five seed questions for each category, covering representative reasoning patterns. Next, we parse both the metadata and seed questions into the LLM agent, which automatically expand them into a larger set of candidate questions. We then conduct a human review process to ensure quality and diversity, where annotators limit the frequency of specific tables, attributes, and operations (e.g., aggregation, area computation, TOP-k queries, “best” selection, or location comparison). Finally, each question is manually rewritten for clarity and naturalness, ensuring that it remains human-readable and provides sufficient semantic hints for both human annotators and LLM agents to generate executable SQL. Next, we describe the procedure used to construct the gold SQL answers in a fair and consistent way.
@@
We employ an LLM agent to generate SQL queries under strict schema and function constraints defined in the metadata. To ensure consistency and interpretability, SQL generation is restricted to a limited yet well-defined set of functions, each aligned with specific reasoning categories. Inspired by TableBench Wu et al. (2025), we incorporate a voting-based validation mechanism in which multiple LLM agents assess the correctness of each generated query. We iteratively refine each question–SQL pair until all three agents reach consensus that the SQL is valid and constitutes a reasonable answer to the question. To further ensure quality, we execute every SQL query against the benchmark database to confirm that it runs successfully and produces a sensible output. Finally, although not illustrated in Figure 1, we perform an additional semantic consistency check in which human reviewers verify that each question faithfully reflects its corresponding SQL logic; any ambiguous or underspecified cases are manually revised for clarity.
4 Experiments
@@ 5 Conclusion
Limitations
This is the section for limitations
Acknowledgments
This is the section for acknowledgments
Appendix A Dataset Details
If there is any discrepancy between the released benchmark and this paper. Please follow the latest released benchmark.
Dataset Rows Points (Before) Points (After) Avg pts (Before) Avg pts (After)
Floodplain 915,760 734,284,445 179,571,368 801.8 196.1
Census tracts 13,444 5,474,058 3,279,789 407.2 244.0
ZCTA 5,284 8,583,835 1,825,650 1624.5 345.5
Table 4: Comparison of geometry statistics before and after simplification. Row counts are preserved while point counts are significantly reduced, improving query efficiency in FloodSQL_Bench.
A.1 Spatial Details
Geometries and preprocessing.
We standardize geometries to OGC:CRS84 (WGS84 lon/lat, degrees). Polygon geometries are retained for floodplain, census_tracts, zcta, and county; claims are stored as point geometries (WKB). In contrast, schools and hospitals are represented by explicit LAT/LON fields (their opaque point BLOBs are discarded), and points are constructed on demand via ST_Point(LON, LAT). Floodplain, tract, ZCTA, and county polygons together contain nearly one million geometries and hundreds of millions of vertices. To ensure query efficiency, we apply topology-preserving simplification (ST_SimplifyPreserveTopology) with dataset-specific tolerances, $10^{-3}$ ($\approx$100m) for floodplains, $10^{-4}$ ($\approx$10m) for tracts, and $5\times 10^{-4}$ ($\approx$50m) for ZCTAs and counties, only on polygons with >>1K–2K vertices, reducing vertex counts by 40–80% without altering query semantics. Table 4 reports this reduction quantitatively.
Point vs. polygon data.
FloodSQL_Bench distinguishes polygon-based from point-based layers. Polygon tables (census_tracts, floodplain, zcta, county) retain geometries for overlay operations. Point-based tables differ: schools and hospitals rely solely on explicit LAT/LON fields, with points constructed on demand (ST_Point(LON, LAT)). Point datasets (claims, schools, hospitals) are consistently treated as explicit latitude/longitude pairs, with geometries constructed on demand. claim geometry is kept in the schema but used only for key-based aggregation via GEOID identifiers. The claims table functions as a key-joinable layer rather than a spatially joined one, enabling direct integration with tract-level indicators (e.g., NRI, SVI, CRE) via 11-digit GEOIDs. This design makes spatial joins explicit and keeps geometry handling transparent for LLMs.
A.2 Value Summary of Tables
As shown in Table 5-14, we provide column statistics and example values for each table in FloodSQL_Bench. For numeric columns we show the value range; for categorical or string columns we show the number of unique values and sample examples. For clarity, we omit raw geometry BLOB columns from the value summary tables. These columns exist in the following datasets: census_tracts, claims, county, floodplain, and zcta.
Column Summary
GEOID 13,444 unique; e.g., 12011010102, 12011110402, 12011010304
STATEFP 3 unique; examples: 12
COUNTYFP 255 unique; examples: 011, 099
NAME 8,048 unique; examples: 101.02, 1104.02
Table 5: Value summary for census_tracts_tx_fl_la.
Column Summary
GEOID 12,756 unique; e.g., 12086000709, 48167723505
STATEFP 3 unique; examples: 12, 22, 48
dateOfLoss 15,042 unique dates
amountPaidOnBuildingClaim range: -201,667.50 .. 10,741,476.93
amountPaidOnContentsClaim range: -80,000.00 .. 500,000.00
amountPaidOnIncreasedCostOfComplianceClaim range: -6,450.00 .. 426,321.00
Table 6: Value summary for claims_tx_fl_la.
Column Summary
GEOID 385 unique; e.g., 48327, 48189
STATEFP 3 unique; examples: 12, 22, 48
COUNTYFP 255 unique; e.g., 327, 189, 011
NAME 357 unique; e.g., Menard, Hale, Armstrong
Table 7: Value summary for county_tx_fl_la.
Column Summary
GFID 199 unique; e.g., 20120202, 20110720LOMR
STATEFP 3 unique; examples: 12, 22, 48
FLD_ZONE 10 unique; e.g., A, AE, X, D
Table 8: Value summary for floodplain_tx_fl_la.
Column Summary
GEOID 5,284 unique; e.g., 34210, 33010, 33919
STATEFP 3 unique; examples: 12, 22, 48
Table 9: Value summary for zcta_tx_fl_la.
Column Summary
HOSPITAL_ID 1,526 unique; e.g., 0005479830, 0019879735
NAME 1,508 unique; e.g., BIG BEND REGIONAL MEDICAL CENTER
ADDRESS 1,511 unique; e.g., 2600 HIGHWAY 118 NORTH
CITY 566 unique; e.g., EL PASO, MIAMI, HOUSTON
STATE 3 unique; TX, FL, LA
ZIP 896 unique; e.g., 79912, 33177
COUNTY 287 unique; e.g., EL PASO, MIAMI-DADE
COUNTYFIPS 306 unique; e.g., 48141, 12086
LAT range: 24.5641986 .. 36.36988585
LON range: -106.5735232 .. -80.05188127
TYPE 10 categories; e.g., GENERAL ACUTE CARE, CRITICAL ACCESS, PSYCHIATRIC
STATEFP 3 unique; 12, 22, 48
UNIQUE_ID 1,526 unique; e.g., 48_hospital_0005479830
Table 10: Value summary for hospitals_tx_fl_la.
Column Summary
SCHOOL_ID 19,669 unique; e.g., 476540, 488332, 226107
NAME 17,947 unique; e.g., UNIVERSITY OF NORTH TEXAS AT DALLAS
ADDRESS 18,299 unique; e.g., 7300 UNIVERSITY HILLS DRIVE
CITY 1,863 unique; e.g., DALLAS, MIAMI, HOUSTON
STATE 3 unique; TX, FL, LA
ZIP 2,874 unique; e.g., 75241, 33156
LAT range: 24.552411 .. 36.49455370700008
LON range: -106.62416140199996 .. -80.03827100000001
TYPE 3 categories; COLLEGE, PUBLIC_SCHOOL, PRIVATE_SCHOOL
STATEFP 3 unique; 12, 22, 48
UNIQUE_ID 19,669 unique; e.g., 48_COLLEGE_476540
Table 11: Value summary for schools_tx_fl_la.
Column Summary
GEOID 13,385 unique; e.g., 12001000201, 12001000202
ST / STATE / ST_ABBR ST range: 12 .. 48; STATE: {Florida, Louisiana, Texas}; ST_ABBR: {FL, LA, TX}
COUNTY 368 unique; e.g., Harris County, Miami-Dade County
FIPS range: 12001000201 .. 48507950302
AREA_SQMI range: 0.0 .. 6,175.6501
E_TOTPOP range: 0 .. 28,999
E_POV150 range: 0 .. 7,002
E_UNEMP range: 0 .. 915
E_AGE65 / E_AGE17 range: 0 .. 9,894 / 0 .. 10,142
E_MINRTY range: 0 .. 21,476
E_MUNIT / E_MOBILE range: 0 .. 6,733 / 0 .. 3,027
EP_* (percentages) many in [-999, 100]; note: -999 denotes missing per SVI convention
RPL_THEMES range: -999 .. 1.0 (percentile ranks with sentinel -999)
Table 12: Selected indicators and ranges for svi_tx_fl_la. (-999 indicates missing values per SVI specification.)
Column Summary
GEOID 13,373 unique; e.g., 12001000201, 12001000202
STATE 3 unique; FL, LA, TX
RFLD_RISKR 7 categories; e.g., Very Low, Relatively Low, Relatively Moderate, Relatively High, Very High, No Rating, Insufficient Data
RFLD_EALB range: 0.0 .. 27,794,853.10
RFLD_EALS range: 0.0 .. 99.99
Table 13: Selected value summary for nri_tx_fl_la.
Column Summary
GEOID 13,444 unique; e.g., 12001000201, 12001000202
GEO_ID 13,444 unique; e.g., 1400000US12001000201
STATE range: 12 .. 48
COUNTY range: 1 .. 507
TRACT range: 100 .. 990200
NAME 13,444 unique; e.g., Census Tract 3.01, Alachua County, Florida
POPUNI range: 0 .. 31,221
PRED0_E range: 0 .. 15,380
PRED0_PE range: 0.0 .. 100.0
PRED12_E range: 0 .. 15,470
PRED12_PE range: 0.0 .. 100.0
PRED3_E range: 0 .. 5,691
PRED3_PE range: 0.0 .. 67.38
Table 14: Selected value summary for cre_tx_fl_la.
A.3 Detailed Join Rules by Table
We provide a per-table summary of how each dataset in FloodSQL_Bench participates in join operations. This complements the join matrix Table 1.
  • county: County-level administrative polygon layer representing U.S. counties. Each record corresponds to a single county area within a state. This table supports both key-based joins via the 5-digit GEOID and spatial joins via the geometry field.
  • census_tracts: Polygon-based spatial layer representing U.S. Census tracts. Each record corresponds to a single tract area defined by the U.S. Census Bureau. This table supports key-based joins via the 11-digit GEOID and spatial joins with polygon-based layers and point-based layers through its geometry field.
  • claims: Key-based table containing National Flood Insurance Program (NFIP) flood claim records. Each record represents a single insurance claim event linked to a census tract. All joins are key-based using the 11-digit GEOID; spatial joins are not applicable.
  • floodplain: FEMA Flood Hazard polygon layer representing physical flood risk zones. Each record corresponds to a FEMA-defined flood hazard area. This table is spatial-only and supports spatial joins through its geometry field, but is not used for key-based joins.
  • zcta: ZIP Code Tabulation Area (ZCTA) polygon layer representing U.S. postal ZIP code regions. ZCTAs are spatial constructs derived from census data and are not identical to USPS ZIP codes. This table supports spatial joins via its geometry field but is not used for key-based joins.
  • schools: Point-based layer representing school facilities across Texas, Florida, and Louisiana. Each record corresponds to one educational institution, including colleges, private, and public schools. This table supports spatial joins through coordinates (LON, LAT) for point-in-polygon operations with polygon layers. ZIP serves as the primary key for key-based joins with hospitals. Geometry can be constructed dynamically using ST_Point(LON, LAT) in WGS84 (EPSG:4326)."
  • hospitals: Point-based layer representing hospital facilities across Texas, Florida, and Louisiana. Each record corresponds to a single healthcare facility. This table supports both spatial joins through coordinates (LON, LAT) and key-based joins with schools via ZIP, with county-level and tract-level datasets via the 5-digit COUNTYFIPS code. Geometry can be constructed dynamically using ST_Point(LON, LAT) in WGS84 (EPSG:4326).
  • SVI: Key-based table representing the Social Vulnerability Index (SVI) dataset published by the CDC/ATSDR. Provides tract-level demographic, socioeconomic, housing, and mobility indicators used to measure a community’s vulnerability to natural hazards and disasters. This dataset contains no geometry field; all joins are key-based via the 11-digit census tract GEOID. A FIPS field identical to GEOID exists in the source data, but GEOID is used as the canonical key in FloodSQL-Bench."
  • CRE: Key-based table representing the Community Resilience Estimates (CRE) dataset published by the U.S. Census Bureau. Provides tract-level measures of population resilience to natural hazards, estimating how well communities can withstand and recover from disasters. This dataset contains no geometry field; all joins are key-based via the 11-digit census tract GEOID.
  • NRI: Key-based table representing the National Risk Index (NRI) dataset developed by FEMA. Provides tract-level estimates of natural hazard risk across the United States, quantifying expected losses, exposure, and composite risk scores for multiple hazard types, including Coastal Flood (CFLD) and Riverine Flood (RFLD). This dataset contains no geometry field; all joins are key-based via the 11-digit census tract GEOID.

### v2

@@ Abstract
Codehttps://github.com/HanzhouLiu/FloodSQL-Bench
Datasetshttps://huggingface.co/datasets/HanzhouLiu/FloodSQL-Bench
Introduction
Text-to-SQL translates natural language questions into executable SQL queries, enabling intuitive access to relational databases through conversational interfaces (Katsogiannis-Meimarakis and Koutrika 2023). General-purpose benchmarks such as Spider (Yu et al. 2018) and BIRD (Li et al. 2023) have significantly advanced research on this task. However, these benchmarks primarily emphasize single-table queries or simple joins in open-domain settings, and thus fail to capture the challenges of domain-specific, multi-table and geospatial reasoning that are crucial for real-world applications.
Flood management is one such domain, where decision-making depends on integrating different data sources, including geographic features, demographic statistics, and critical infrastructure records (Cutter, Boruff, and Shirley 2012; Wing et al. 2018). Unlike open-domain settings, practitioners must combine information across multiple layers, for example linking census tracts with floodplain boundaries or associating hospitals with historical flood claims. To overcome these challenges, we propose FloodSQL-Bench, a novel benchmark constructed from real-world datasets and designed around progressively complex queries, ranging from single-table lookups to multi-table reasoning involving key-based joins, spatial joins, and their hybrids. This design enables systematic evaluation of Text-to-SQL models on realistic geospatial and socio-economic reasoning tasks essential for flood risk assessment and emergency response.
@@ Introduction
Text-to-SQL research has evolved from single-table parsing to cross-domain and multi-table reasoning. Early datasets such as ATIS and GeoQuery focused on domain-specific grammars, whereas Spider (Yu et al. 2018) established a large-scale cross-domain benchmark that has become the de facto standard for evaluating Text-to-SQL systems. Subsequent works, including SParC (Yu et al. 2019b) and CoSQL (Yu et al. 2019a), extended Spider to multi-turn and conversational settings. More recently, benchmark efforts such as TableBench (Wu et al. 2025) target broader table-based reasoning beyond SQL generation. However, all of these benchmarks primarily center on open-domain databases with relatively simple join structures, and rarely address domain-specific, multi-table reasoning that involves heterogeneous relational and spatial data, a key challenge for high-stakes applications such as flood risk management.
@@
Geospatial reasoning introduces unique challenges beyond conventional relational semantics, as queries often involve spatial joins, coordinate transformations, and topology-aware aggregation. Prior work on natural-language interfaces to spatial databases (Li et al. 2019; Liu et al. 2025) and spatial question answering (Chen et al. 2021) has explored spatial relations such as containment, intersection, and proximity. However, these systems are typically evaluated on synthetic or small-scale datasets and do not address realistic, domain-specific reasoning that integrates heterogeneous spatial and tabular information. In contrast, real-world flood management requires the fusion of diverse geospatial sources—such as physical floodplains, demographic indicators, and critical infrastructure layers, to support actionable decision-making (Cutter, Boruff, and Shirley 2012; Wing et al. 2018). Despite their importance, no existing Text-to-SQL or QA benchmarks systematically incorporate such spatially grounded, multi-table reasoning tasks. FloodSQL-Bench bridges this gap by providing a unified benchmark that combines key-based, spatial, and hybrid joins across real disaster-management datasets, enabling rigorous evaluation of large language models (LLMs) under retrieval-augmented (RAG) settings.
@@
Retrieval-Augmented Generation (RAG) has become a central approach for grounding large language models (LLMs) in traditional NLP tasks like question answering and conversation (Lewis et al. 2020; Izacard and Grave 2021; Shuster et al. 2021), with recent work extending retrieval beyond unstructured text to structured sources such as relational databases and tables (Ayala and Bechard 2024; Wu et al. 2024). However, existing text-to-SQL benchmarks primarily focus on non-spatial schema and do not evaluate reasoning over geometric data, spatial joins, or multi-layer geospatial infrastructures. Although prior work has explored geospatial question answering using knowledge graphs (Li et al. 2025) and GIS tool–driven workflow automation (Zhang et al. 2024), neither line of research concentrates on SQL-based spatial joins across multiple tables. Recent advances in table-aware retrieval (Zhang et al. 2023; Ziletti and DAmbrosi 2024) further highlight the importance of retrieval granularity for structured reasoning, yet these methods have not been explicitly tested in spatial settings where geometry, topology, and coordinate systems must be considered. To the best of our knowledge, FloodSQL-Bench addresses these gaps by providing the first benchmark designed for multi-layer geospatial SQL reasoning and RAG evaluation, enabling systematic assessment of LLMs across lookup, relational, and spatially grounded analytical queries.
Benchmark Construction
@@
To balance realism with tractability, FloodSQL-Bench focuses on three flood-prone states, Texas, Florida, and Louisiana, which together account for a disproportionate share of National Flood Insurance Program (NFIP) claims and Federal Emergency Management Agency (FEMA)–declared disasters. Within this scope, FloodSQL-Bench integrates ten heterogeneous tables across three spatial aspects: (i) non-spatial layers, claims, svi111svi: Social Vulnerability Index., cre222cre: Community Resilience Estimates., nri333nri: National Risk Index.; (ii) polygon layers, floodplain, census_tracts, zcta, county; and (iii) point layers, schools, hospitals. These layers reflect key information needed for flood risk management (Cutter, Boruff, and Shirley 2012; Wing et al. 2018).
@@
Tabular Data Foundations
@@
FloodSQL-Bench adopts standardized geographic identifiers defined by the U.S. Census Bureau to ensure consistent key-based joins across heterogeneous tables (U.S. Census Bureau 2020). At the tract level, an 11-digit GEOID encodes a 2-digit state code, a 3-digit county code, and a 6-digit tract code, serving as the primary join key among the census_tracts, claims, svi, nri, and cre tables. At the county level, the 5-digit prefix of the tract GEOID uniquely identifies each county in the county and hospitals tables. At the ZIP level, the ZIP field serves as the geographic identifier for the schools and hospitals tables. Two spatial layers, zcta and floodplain, lack non-spatial join keys and are instead linked through spatial relationships only.
@@
FloodSQL-Bench standardizes spatial representations across various data sources. Specifically, polygon layers retain geometries to support polygon–polygon joins via spatial SQL functions such as ST_Intersects, ST_Contains, and etc. In contrast, point layers store only explicit latitude and longitude (LAT&LON) fields, with geometries constructed on demand using ST_Point(LON, LAT). To ensure efficiency, all geometries are projected to a common coordinate reference system (CRS) (Goodchild 1992; Burrough, McDonnell, and Lloyd 2015) and simplified with topology-preserving tolerances (Visvalingam and Whyatt 2017; GEOS Development Team ).
Join Semantics and Relations
@@
We follow the procedure of TableBench (Wu et al. 2025) to construct diverse question–SQL pairs. Specifically, we first manually design five seed questions for each category, covering representative reasoning patterns. Next, we parse both the metadata and seed questions into the LLM agent, which automatically expand them into a larger set of candidate questions. We then conduct a human review process to ensure quality and diversity, where annotators limit the frequency of specific tables, attributes, and operations (e.g., aggregation, area computation, TOP-k queries, “best” selection, or location comparison). Finally, each question is manually rewritten for clarity and naturalness, ensuring that it remains human-readable and provides sufficient semantic hints for both human annotators and LLM agents to generate executable SQL. Next, we describe the procedure used to construct the gold SQL answers in a fair and consistent way.
@@
We employ an LLM agent to generate SQL queries under strict schema and function constraints defined in the metadata. To ensure consistency and interpretability, SQL generation is restricted to a limited yet well-defined set of functions, each aligned with specific reasoning categories. Inspired by TableBench (Wu et al. 2025), we incorporate a voting-based validation mechanism in which multiple LLM agents assess the correctness of each generated query. We iteratively refine each question–SQL pair until all three agents reach consensus that the SQL is valid and constitutes a reasonable answer to the question. To further ensure quality, we execute every SQL query against the benchmark database to confirm that it runs successfully and produces a sensible output. Finally, although not illustrated in Figure 1, we perform an additional semantic consistency check in which human reviewers verify that each question faithfully reflects its corresponding SQL logic; any ambiguous or underspecified cases are manually revised for clarity.
Experiments
@@ Conclusion
Acknowledgment
This work used the DeltaAI system at the National Center for Supercomputing Applications [award OAC 2320345] through allocation [allocation number CIV250031] from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.